OpenStudy (anonymous):

what is a variable range in algebra 1A?

OpenStudy (anonymous):

The range is the extremeties of the y coordinate.

OpenStudy (anonymous):

extremity*

OpenStudy (anonymous):

so how would i find that in a distance=time X speed problem?

OpenStudy (anonymous):

An option is to graph it on Geogebra.

OpenStudy (anonymous):

if a distance is 480 miles and the starting point is zero would it be 0 to 480=variable range?

OpenStudy (anonymous):

And comparing it to a real life situation, the distance can not be negative.

OpenStudy (anonymous):

Yes, that is correct.

OpenStudy (anonymous):

so that is correct then? that would be my answer?

OpenStudy (anonymous):

There's a specific way how to write your answer.

OpenStudy (anonymous):

can you tell me?

OpenStudy (anonymous):

\[\left\{ y \epsilon \mathbb{R}, 0 \le d \le 480 \right\}\]

OpenStudy (anonymous):

aha ok thank you @GodlyMaths